OpenStudy (anonymous):

let T be multiplication by the matrix A= (top: 1,0 -2; bottom: -2,1,2) a) find a basis for the range of T b) find a basis for the kernel of T c) the rank and nullity of T d) the rank and nullity of A Please, explain me.

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 0 & -2 \\ -2 & 1 & 2\end{matrix}\right]\]

OpenStudy (anonymous):

Have you row reduced the matrix yet? All of the answers to your questions can be found by first row reducing the matrix.

OpenStudy (anonymous):

its rref is \[\left[\begin{matrix}1 & 0 & -2\\ 0 & 1 & -2\end{matrix}\right]\]

OpenStudy (anonymous):

Alright. Lets knock out the easy ones first. How many pivot variables does that matrix have? and how many free variables are there? (I'm not sure what your class is calling them, so if you dont know the definitions of some terms I say, let me know).

OpenStudy (anonymous):

2 pivots, and 0 free variable

OpenStudy (anonymous):

You are correct about the pivots, but there is actually one free variable. It goes by the columns. So two columns have leading ones in them (those are the pivots), and one column doesnt have a leading one (the last one).

OpenStudy (anonymous):

oh, yea, my bad, 1 free

OpenStudy (anonymous):

but my logic is 2 equations, 3 variables, ---> 1 free

OpenStudy (anonymous):

The rank of A = number of pivots The nullity of A = number free variables. The rank and nullity of T are the same as the rank and nullity of A (since A is the matrix representation of T).

OpenStudy (anonymous):

so the rank of T is 2 nullity of A is 1?

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

Hi friend, I confused. It sounds like T =A?

OpenStudy (usukidoll):

I don't think so...different meanings, but then again I have yet to reach that section

OpenStudy (anonymous):

@UsukiDoll LOL,

OpenStudy (anonymous):

T is a linear transformation A is simply a matrix

OpenStudy (anonymous):

T(x)= Ax, so whether you are talking about a transformation (T), or matrix multiplication (A), they mean the same thing. They are linked.

OpenStudy (anonymous):

while they go hand-in-hand, A is just one of the many possible matrix representations of T.

OpenStudy (usukidoll):

oh man... x*( like I said I'm not on that sectyion

OpenStudy (usukidoll):

well now I know something in advance ;)

OpenStudy (usukidoll):

This is Linear Algebra stuff right?

OpenStudy (anonymous):

who needs matrices anyway

OpenStudy (anonymous):

yes, this is Linear Algebra.

OpenStudy (usukidoll):

ok I'm on the same page except I'm struggling with a proof and need a refresher course on subspaces...the ones with the u+v and cu properties.

OpenStudy (anonymous):

@joemath314159 so, in simple way, I don't get. more explanation please :)

OpenStudy (anonymous):

what don't you get Hoa?

OpenStudy (usukidoll):

everything =)

OpenStudy (anonymous):

do not underestimate Hoa, Usukidoll :)

OpenStudy (anonymous):

whether A =T or not?

OpenStudy (anonymous):

oh. that is comparing apples and oranges, Hoa. they are related, but not the same thing A is a matrix T is a transformation which COULD be represented by A.

OpenStudy (anonymous):

got it, continue the stuff, I want a and b parts

OpenStudy (usukidoll):

umm I think he went

OpenStudy (anonymous):

probably need @joemath314159 for that. i don't know how to write matrices

OpenStudy (usukidoll):

wahhh I need my proof correction to be checked too. ;__:

OpenStudy (anonymous):

we need someone really good with matrices~ i'm not

OpenStudy (anonymous):

so, use my way, (top..., bottom...)

OpenStudy (usukidoll):

this better be a connection error. I'm gonna be so sad if it's not

OpenStudy (usukidoll):

or draw the matrix.

OpenStudy (e.mccormick):

@TaylorNeedsHelp What, write it in latex? Use begin matrix and seperate the parts with & and use \\ for new lines.

OpenStudy (usukidoll):

drawing is easier

OpenStudy (anonymous):

what's latex?

OpenStudy (usukidoll):

does anyone know proof writing

OpenStudy (usukidoll):

program for writing math formulas

OpenStudy (anonymous):

ok, you can go to Latex Practicing by clicking Find More Subjects and practice your skill there

OpenStudy (anonymous):

hehehe.... good luck!!!

OpenStudy (anonymous):

in time.

OpenStudy (e.mccormick):

@TaylorNeedsHelp LaTeX is a formatting language tha got adapted into the MathML thing used to show equations here.

OpenStudy (usukidoll):

O___O

OpenStudy (anonymous):

that's how the people are writing those fancy equations ?

OpenStudy (anonymous):

or you can use those post to practice right here, I am ok.

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

hey, UsukiDoll, are you good with matrices? :)

OpenStudy (usukidoll):

if you know someone who can help me correct a proof...

OpenStudy (usukidoll):

gawd I don't even know if my corrections is in the right track or went to spaceville

OpenStudy (anonymous):

tit for tat, UsukiDoll? Show me your proof

OpenStudy (usukidoll):

I'm on I think...basis and dimensions for the matrices

OpenStudy (e.mccormick):

The Kernel is the solution set, or zeros of it.

OpenStudy (usukidoll):

but I don't know subspaces and subsets too good.

OpenStudy (usukidoll):

that's why I'm here lol

OpenStudy (anonymous):

you're in luck, UsukiDoll, I'm ok with vector spaces

OpenStudy (usukidoll):

ok do you know how to prove whether or not the tests for u+v and cu work?

OpenStudy (usukidoll):

or do I just plug stuff into the equations/axioms?

OpenStudy (anonymous):

post your question.

OpenStudy (anonymous):

not that i have to tell you that

OpenStudy (usukidoll):

can't I'm stuck with a proof question

OpenStudy (e.mccormick):

So the basis of the kernel is a basis that is related to the parameteric equation that solves for the null space of A.

OpenStudy (usukidoll):

unless you want me to post here.

OpenStudy (anonymous):

get Hoa's permission

OpenStudy (usukidoll):

Hoa can I hijack your thread? >:)

OpenStudy (anonymous):

do threads usually get this long?

OpenStudy (e.mccormick):

The range, well, that is the output of the transformation. Hmm. Not sure how to get the range of the results.

OpenStudy (usukidoll):

my own thread is lasting for three days and counting

OpenStudy (anonymous):

what thread? link

OpenStudy (e.mccormick):

Or basis of the results... I mean.

OpenStudy (usukidoll):

like one of the questions would be Which are the following are subsets? The set of all nxn symmetric matrices The set of all nxn diagonal matrices The set of all nxn nonsingular matrices do I just use the axiom formulas to determine if there is such a thing or?

OpenStudy (anonymous):

subsets need a superset

OpenStudy (usukidoll):

what the!

OpenStudy (anonymous):

You cannot define subsets if they are not subsets of anything... right, @Hoa ?

OpenStudy (usukidoll):

what is a superset?

OpenStudy (anonymous):

all you go to far and go around. since it fromR^3--> R^2 so, the rank(t) =2 and the basis of R^2 is (0,1) and (1,0) hehehe...It sounds right?

OpenStudy (usukidoll):

what the heck?

OpenStudy (usukidoll):

what? I said heck not a swear word

OpenStudy (anonymous):

to me, all of UsukiDoll ask for about the subsets are under the "Square Matrices"

OpenStudy (anonymous):

Set

OpenStudy (anonymous):

hihi.... sorry for breaking the sentence down.

OpenStudy (anonymous):

I am not good at matrices, that is why i am here to make question, the good thing is I understand what I learn hihihi.... all you make question, If I know, I will answer hihi..

OpenStudy (e.mccormick):

My prof has been out 3 days this semester, itherwise I would be far enough a long to perhaos have more help. This stuff is coming up soon in our class. Have not looked ahead much because doing test prep.

OpenStudy (anonymous):

no question? friends? if so, I gonna close the question

OpenStudy (anonymous):

k, I think your your basis for the kernel of T is going to be \[\left(\begin{matrix}2 \\ 2\\1\end{matrix}\right)\] and the basis for the range of T is the two pivot columns of the original matrix \[ \left(\begin{matrix}1 \\ -2\end{matrix}\right) and \left(\begin{matrix}0 \\ 1\end{matrix}\right)\]

OpenStudy (anonymous):

If you want to see a reference video, this is what I was looking at: http://www.youtube.com/watch?v=Vmr6EAQQazA

Directrix (directrix):

@Hoa Some good reading here: http://tinyurl.com/cxr5rgo